### v1

@@ 3.3 Spatial-Aware Block-Wise Feature Aggregator
To implement $g_{\varphi}$ in Eq. 5, SABW applies a ResNet [15] layer to introduce timesteps by adding the time embedding into the feature map, and a Transformer [54] layer to introduce textual prompts by the cross-attention,
@@
3.4 Training
Given the ensembled denoiser with AFA $\mathcal{F}_{\Theta}$, the training object is the same as the denoising loss in Eq. 1,
$\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0},\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{c},t}\left\|\mathcal{F}_{\Theta}\left(\mathbf{x}_{t},\mathbf{c},t\right)-\boldsymbol{\epsilon}\right\|^{2}\ .$ (12)
@@ 3.4 Training
Table 1: Quantitative comparison between our AFA with the three base models in Group I (i.e., ER, MMR, and RV) and the baseline methods.
COCO 2017 Draw Bench Prompts
FID $\downarrow$ IS CLIP-I CLIP-T AES PS HPSv2 IR
Base Model A 13.01 5.65 .6724 .2609 5.4102 21.6279 27.8007 .3544
Base Model B 13.45 5.43 .6775 .2652 5.5013 21.4624 27.7246 .2835
Base Model C 12.32 6.32 .6890 .2566 5.4881 21.8031 27.9652 .3922
Weighted Merging [56] 10.65 6.93 .6861 .2626 5.4815 21.7272 27.9086 .3909
MBW [33] 11.03 6.51 .6870 .2624 5.4812 21.7201 27.9080 .3922
autoMBW [4] 13.35 5.51 .6772 .2577 5.5056 21.4785 27.8192 .3672
MagicFusion [68] 10.53 6.85 .6751 .2620 5.3431 21.3840 27.8105 .3317
AFA (Ours) 9.76 7.14 .6926 .2675 5.5201 21.8263 27.9734 .4388
3.5 Inference
The inference of AFA starts from the sampled Gaussian noise. Then, the diffusion scheduler (e.g., DDIM [51], PNDM [27], DPM-Solver [29, 30], etc.) is applied to generate images with multiple denoising steps.
For each inference step, the noise prediction relies on the technique of Classifier-Free Guidance (CFG) [18], which is formulated as
$\hat{\boldsymbol{\epsilon}}_{t}^{\text{pred}}=\hat{\boldsymbol{\epsilon}}_{t}^{\text{uc}}+\beta_{\text{CFG}}(\hat{\boldsymbol{\epsilon}}_{t}^{\text{c}}-\hat{\boldsymbol{\epsilon}}_{t}^{\text{uc}})\ .$ (13)
Among them, $\hat{\boldsymbol{\epsilon}}_{t}^{\text{pred}}$, $\hat{\boldsymbol{\epsilon}}_{t}^{\text{uc}}$, $\hat{\boldsymbol{\epsilon}}_{t}^{\text{c}}$, and $\beta_{\text{CFG}}$ are the final predicted noise, the predicted noise with condition, the predicted noise without condition, and a guidance scale, respectively. The latent for the next step is
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\mathbf{x}_{t}-\frac{\sqrt{1-\alpha_{t}}}{\sqrt{\alpha_{t}}}\hat{\boldsymbol{\epsilon}}_{t}^{\text{pred}}\ .$ (14)
Finally, the generated image is achieved from the latent by a VAE decoder.
4 Experiments
4.1 Experiment Settings
4.1.1 Base Models.
We select six popular models with the same architecture from SDv1.5 [49] in CivitAI. There are twelve down-sampling blocks, one middle block, and twelve up-sampling blocks for each model. These models are randomly divided into two groups, with three models in each group. The models from the same group will be ensembled.
  • Group I contains EpicRealism (ER) [9], MajicMixRealistic (MMR) [31], and RealisticVision (RV) [42];
  • Group II contains AbsoluteReality (AR) [1], CyberRealistic (CR) [6], and RealCartoonReslistic (RCR) [41].
Table 2: Quantitative comparison between our AFA with the three base models in Group II (i.e., AR, CR, and RCR) and the baseline methods.
COCO 2017 Draw Bench Prompts
FID $\downarrow$ IS CLIP-I CLIP-T AES PS HPSv2 IR
Base Model A 12.12 5.66 .6849 .2623 5.5641 21.8013 28.0183 .4238
Base Model B 12.41 5.59 .6818 .2580 5.5027 21.7249 28.0343 .4202
Base Model C 12.05 5.95 .6638 .2642 5.5712 21.4936 27.8089 .3367
Weighted Merging [56] 11.53 6.56 .6824 .2631 5.5756 21.7516 28.0014 .4387
MBW [33] 12.06 6.42 .6826 .2632 5.5772 21.7487 28.0029 .4396
autoMBW [4] 12.39 5.62 .6774 .2588 5.5478 21.5135 27.9873 .3513
MagicFusion [68] 11.63 7.13 .6790 .2640 5.4674 21.4270 27.9608 .4194
AFA (Ours) 10.27 7.42 .6855 .2717 5.5798 21.8059 28.0371 .4892
4.1.2 Training.
AdamW [28] is used as the optimizer with a learning rate of 0.0001 and a weight decay of 0.01. Note that only the parameters of SABW are optimized, while those of U-Net denoisers are frozen. Our AFA framework is trained on 10,000 samples from the dataset JourneyDB [37] for 10 epochs with batch size 8. To enable CFG, we use a probability of 0.1 to drop textual prompts.
4.1.3 Evaluation Protocols.
For a fair comparison, all the methods generate 4 images by DDIM [17] for 50 inference steps. The CFG weight $\beta_{\text{CFG}}$ in Eq. 13 is set to 7.5. We evaluate the models with two datasets, which are COCO 2017 [26] and Draw Bench Prompts [48], respectively.
  • COCO 2017 comprises 118,287 and 5,000 image-caption pairs in the test and validation sets. All the models generate images with a resolution of 256$\times$256. We apply four metrics to evaluate the generation performance, which are Fréchet Inception Distance (FID), Inception Score (IS), CLIP-I, and CLIP-T, respectively. FID and IS are applied to the test set, while CLIP-I and CLIP-T are applied to the validation set. Both FID and IS are metrics to assess the quality of generated images. Note that lower FID means better quality. CLIP-I is the similarity between the CLIP [38] images embeddings of generated images and that of images from the image-caption pairs. CLIP-T is the CLIPScore [16] between the generated images with the captions.
  • Draw Bench Prompts contains 200 evaluation prompts. All the models generate images with a resolution of 512$\times$512. We apply 4 evaluation metrics, which are AES [2], Pick Score (PS) [23], HPSv2 [59], and Image Reward (IR) [60], respectively. All three metrics evaluate the performance by a model that simulates human preferences. The evaluations are conducted 20 times to ensure statistical significance.
4.1.4 Baselines.
We compare our AFA with several methods, including three merging-based methods (i.e., Weighted Merging [56], MBW [33], and autoMBW [4]) and one ensembling-based method (i.e., MagicFusion [68]). These baselines are also evaluated on two evaluation protocols under two model groups.
Figure 4: Effect of varying inference steps.
4.2 Quantitative Comparison
The results of the two evaluation protocols are summarized in Tab. 1 and Tab. 2. Generally, compared with base models, performance improvements are achieved by our AFA. For COCO 2017, after being ensembled by AFA, FID, IS, CLIP-I, and CLIP-T exhibit enhancements under both Group I and Group II. This indicates that the quality of our generated images is more closely with the dataset images. Additionally, the contexts of our generated images align more closely with those of the dataset images and corresponding captions. For Draw Bench Prompts, all four metrics show improvements under both Group I and Group II. The improvement on AES suggests that the aesthetics of our generated images surpass those generated by the base models. The improvements on PS, HPSv2, and IR suggest that the general generation capability can be boosted by applying AFA. Overall, the quantitative results validate the effectiveness of our AFA in ensembling text-to-image diffusion models.
As depicted in Fig. 4, we use AFA to ensemble two models from Group I or Group II, and present the IR scores of base models and the ensembled models on Draw Bench Prompts. The results show that AFA significantly improves the performance when ensembling two U-Net denoisers. The results presented in Tab. 1, Tab. 2, and Fig. 4 indicate that AFA can be effectively applied to various model combinations and can be extended to accommodate different numbers of models. Simultaneously, it consistently achieves superior generation performance compared to a single base model.
4.3 Qualitative Comparison
As shown in Fig. 5, we present a qualitative comparison of the results from the base models, some of the baseline methods, and our AFA, which allows us to make several insightful observations. Firstly, AFA can generate images with better aesthetics. For example, in Fig. 5 (a) and (c), AFA can generate images with improved composition and finer details, compared to both the base models and the baseline methods. Secondly, AFA excels in achieving superior context alignment. For example, in Fig. 5 (b), all the base models and the baseline methods generate images containing more than one clock, despite the textual prompt specifying a single clock. In contrast, only our AFA generates the image with just one clock, accurately reflecting the provided context. Thirdly, AFA can focus on the correct context of the base models and drop out the incorrect context. For example, in Fig. 5 (d), only the base model B generates the image with two dogs, while the baseline methods do not fully trust the base model B, and generate the images with people. Only our AFA fully trusts the base model B, and generates the image with two dogs, which aligns with the textual context.
Figure 6: Visualization of the learned attention maps. In each attention map, the position with a lighter color means a higher attention weight. Conversely, the position with a darker color means a lower attention weight.
4.4 Model Study
4.4.1 Visualization of Attention Maps.
As shown in Fig. 6, we select to visualize the attention maps of the first and the penultimate (i.e., 24-th) blocks, both of which are learned by ensembling models in Group II through our AFA. For the first block, the learned attention maps seem to concentrate solely on the feature map of AR at the first inference stage. However, at the last few inference steps (i.e., from the 30th to the 50th step), the roles of the feature maps from different models begin to diverge. The attention maps tend to highlight the contextual body (i.e., the astronaut and the horse in Fig. 6 (a), and the pizza and the oven in Fig. 6 (b)) in the feature map of RCR, while focusing on the background elements in the feature maps of AR and CR. For the penultimate block, as the inference progresses, the learned attention maps gradually emphasize the contextual body within the feature map of AR, while concentrating on the background within the feature maps of both CR and RCR. Overall, the visualization of the learned attention maps indicates that our AFA can effectively ensemble diffusion models based on the context and the timesteps.
4.4.2 Effect of Fewer Inference Steps.
As shown in Fig. 4 and Fig. 7, we investigate the resilience of our AFA to a reduction in the inference steps. Fig. 4 clearly shows a significant drop in the performance of the base models when the number of inference steps is reduced. However, the performance of our AFA remains almost stable when the number of inference steps is reduced from 50 to 20, and only experiences a slight decline when the number of inference steps is reduced to 10. Additionally, in Fig. 7, we can find that the quality of the images generated by the base models deteriorates with the reduction of inference steps. In contrast, our AFA stably generates high-quality images. Overall, it indicates that greater tolerance for fewer inference steps will be achieved after ensembling by AFA. Furthermore, while a single inference step may require a proportional increase in time, the total time consumed by the entire inference process does not escalate in the same manner.
Figure 7: Qualitative comparison under different inference steps.
4.4.3 Ablation Study.
As shown in Tab. 3, we conduct the ablation study to investigate the necessity of each component within our AFA. Firstly, we conduct experiments with the following settings. (1) Only Ensembling Last Block: The AFA ensembles only the last block, leaving all other blocks unaltered. (2) Block-Wise Averaging: Each block is ensembled by averaging the output feature maps, rather than using SABW. (3) Noise Averaging: Only the last block is ensembled by averaging the output feature maps (i.e., predicted noises). The observed performance declines in these settings suggest that the design of the AFA effectively contributes to the ensembling of diffusion models, thereby leading to improved generation. Next, we conduct an ablation study on the spatial location from AFA, which learn only a single attention score (i.e., a scalar) for the feature of each U-Net denoiser, as opposed to an attention map. The observed decrease in performance suggests that learning a spatial attention map to adaptively adjust the weights of various U-Net denoisers in the level of spatial location is beneficial for the ensemble of different diffusion models. Finally, we individually ablate the timestep and the textual condition which are introduced to AFA. The decreasing performance indicates that both the introduced timestep and textual condition play crucial roles in ensembling diffusion models by AFA.
4.4.4 Discussion on Computational Efficiency.
(1) Ensembling Methods. Compared with another ensembling method (i.e., MagicFusion), the computational efficiency of our AFA for a single inference step aligns closely with it. Both AFA and MagicFusion necessitate running all base models in one inference step. Although SABW in AFA introduces additional parameters, which leads to more computations, it has fewer than the base models (about 1/16). To be precise, our AFA may be marginally less efficient than MagicFusion on single inference steps, but the difference is practically negligible. (2) Merging Methods. Because the various models are collapsed into one model, the computational efficiency of the merging methods equals that of a single base model. For the base models and the merging methods, a single image generation takes about 7 seconds with 50 inference steps. When ensembling three such base models, a single generation will take about 22 seconds with the same inference steps. Compared with ensembling, the computation efficiency of one inference step of the merging methods is significantly higher. However, as we discussed in Effect of Fewer Inference Steps of Sec. 4.4, AFA shows remarkable tolerance for fewer inference steps. When limited to 20 inference steps, the performance of AFA does not significantly decline (as shown in Fig. 4), while the time consumed is reduced to about 8.8 seconds. This suggests that AFA has comparable computational efficiency with that of the base models or the merging methods throughout the entire inference process.
5 Conclusion
Table 3: Ablation study of AFA.
IR (Group I) IR (Group II)
AFA (Full Model) .4388 .4892
Only Ensembling Last Block .4176 .4374
Block-Wise Averaging .4001 .4372
Noise Averaging .3919 .4355
w/o Spatial Location .4044 .4610
w/o Timestep .4235 .4622
w/o Textual Condition .4163 .4559
In this paper, we aim to ensemble multiple diffusion models to achieve better generation performance. To this end, we propose the Adaptive Feature Aggregation (AFA) method, which dynamically adjusts the contributions of multiple models at the feature level according to different prompts, initial noises, denoising steps, and spatial locations. Specifically, we design a lightweight Spatial-Aware Block-Wise feature aggregator that produces attention maps to adaptively aggregate the block-level intermediate features from multiple U-Net denoisers. Both the quantitative and qualitative experiments demonstrate that AFA achieves improvements in image quality and context alignment.
5.0.1 Limitation.
A significant limitation of our AFA is that a single inference step may necessitate a proportional increase in inference time, resulting in lower computational efficiency for that step. Fortunately, due to its high tolerance for fewer inference steps, the total time consumed by the entire inference process does not escalate in the same manner. This results in a computational efficiency that is on par with that of the base models or the merging methods.

### v2

@@ Appendix F Ensembling by Mixture-of-Experts
To implement the router $r_{\zeta}^{(j)}$, we apply a ResNet layer to introduce timesteps by adding the time embedding into the feature, and a Transformer layer to introduce textual prompts by the cross-attention. Specifically, given the input for the $j$-th block (i.e., $\mathbf{x}_{t}^{(j)}\in\mathbb{R}^{h_{j}\times w_{j}\times c_{j}}$), the output logits for $N$ models (i.e,, $\mathbf{l}_{t}^{(j)}$) are achieved by
$\mathbf{h}_{t}^{(j)}$ $\leftarrow\text{ResLayer}(\mathbf{x}_{t}^{(j)}+\gamma(t))\in\mathbb{R}^{h_{j}\times w_{j}\times d}\ ,$ (18)
@@
$\mathbf{l}_{t}^{(j)}$ $\leftarrow\text{Linear}(\text{AvgPool}(\mathbf{o}_{t}^{(j)}))\in\mathbb{R}^{N}\ .$ (20)
Although the time required for a single forward pass in the MoE methods is approximately equivalent to that of a single model, they face challenges in performing parallel computations when generating multiple images simultaneously. This limitation becomes evident during classifier-free guidance. If the router selects two different models to process the conditioned and unconditioned inputs, MoE can only execute two sequential forward passes, rather than concurrently processing these two scenarios within a single forward pass. Moreover, when generating $K$ images at once, the MoE method has a high probability of necessitating $2K$ forward passes. While these $2K$ inputs can be divided among $N$ models, this only reduces the complexity associated with the number of forward passes from $O(K)$ to $O(N)$. This reduction still represents a substantial computational demand.
The quantitative comparison among the base models, the MoE methods, and our AFA is summarized in Table 4. AFA demonstrates superior performance compared to the MoE methods. This may be because the base denoisers have well-balanced denoising capabilities across all timesteps, and selecting a specific denoiser for a given timestep does not significantly enhance the overall denoising capability of the ensembled model. Furthermore, we compare AFA and the MoE methods using fewer inference steps (i.e., 20 steps). As shown in Table 5, unlike our AFA, the MoE methods do not demonstrate robustness with fewer inference steps, performing similarity to the base models.
The efficiency comparison among the base models, the MoE methods, and our AFA is summarized in Table 6. Firstly, the MoE methods have a similar number of parameters and training TFLOPs as our AFA. This is because the MoE methods still require all base model to be available for selection, and during training, each base model must perform a forward pass. Secondly, during inference, the MoE methods exhibit lower TFLOPs compared to AFA, as only one base model is activated at each timestep. However, despite this advantage in TFLOPs, the inference time of the MoE methods is comparable to that of AFA due to their inability to leverage parallel inference effectively. Moreover, the MoE methods do not support fewer inference steps. In contrast, when AFA reduces the number of inference steps to 20, it still outperforms the MoE methods, even when they use 50 inference steps, while also achieving lower TFLOPs and faster inference. This highlights the efficiency and effectiveness of our AFA.
Group I Group II
AES PS HPSv2 IR AES PS HPSv2 IR
Base Model A 5.4102 21.6279 27.8007 .3544 5.5641 21.8013 28.0183 .4238
Base Model B 5.5013 21.4624 27.7246 .2835 5.5027 21.7249 28.0343 .4202
Base Model C 5.4881 21.8031 27.9652 .3922 5.5712 21.4936 27.8089 .3367
Denoiser-Level MoE 5.4727 21.6893 27.5345 .4019 5.5638 21.5531 27.7843 .4380
Block-Level MoE 5.5037 21.7690 27.8749 .4037 5.5718 21.7636 27.9842 .4463
AFA (Ours) 5.5201 21.8263 27.9734 .4388 5.5798 21.8059 28.0371 .4892
@@
Table 4: Quantitative comparison between AFA with the base models and the MoE methods.
Group I Group II
AES PS HPSv2 IR AES PS HPSv2 IR
Denoiser-Level MoE 1.2310 2.9435 4.1739 .0184 2.4556 3.6723 5.3415 .0237
Block-Level MoE 1.6854 3.4573 5.3583 .0098 2.5734 4.2368 4.7252 .0344
AFA (Ours) 5.4951 21.8093 27.7347 .4191 5.5322 21.7830 27.9775 .4803
Table 5: Quantitative comparison of AFA and the MoE methods with 20 inference steps.
1 image 2 images 4 images
# params. T-TLOPs I-TFLOPs Times (s) I-TFLOPs Times (s) I-TFLOPs Times (s)
Base Model 859.52M 70.24 2.94 140.45 5.61 280.88 6.87
Denoiser-Level MoE 2581.01M 6.36 72.37 5.36 144.75 10.37 289.48 20.81
Block-Level MoE 2621.03M 6.43 74.41 5.39 148.82 10.43 297.65 21.00
AFA (50 inf. steps) 2621.03M 6.42 218.74 8.98 437.48 15.57 874.96 21.33
AFA (20 inf. steps) 2621.03M 6.42 86.61 3.62 173.22 6.41 346.45 8.74
Table 6: Efficiency Comparison between AFA with the base models and the MoE methods when ensembling three base models to generate images with a resolution of 512$\times$512. T-TFLOPs and I-TFLOPs denote TFLOPs for training and inference, respectively. Note that only the parameters of denoisers are considered. Note that all the inference times are evaluated in our A6000 environment over 20 runs. Note that both the base models and MoE methods are evaluated with 50 inference steps, because they cannot achieve comparable performance with fewer inference steps.
Appendix G Ensembling More Models
To evaluate the scalability of our AFA, we apply it to ensemble a larger number of base models. Specifically, we select and ensemble several base models with best performance, and test IR metrics in Draw Bench Prompts. As shown in Table 7, ensembling more base models leads more performance improvements. Additionally, increasing the number of the ensembled base models enhances the tolerance for fewer inference steps.
1 base model 2 base models 3 base models 4 base models 5 base models 6 base models
50 inference steps 0.4238 0.5003 0.4892 0.4967 0.5042 0.5347
40 inference steps 0.4065 0.4834 0.4874 0.4906 0.4984 0.5310
30 inference steps 0.3593 0.4791 0.4846 0.4853 0.4993 0.5294
20 inference steps 0.2103 0.3641 0.4803 0.4840 0.4975 0.5304
10 inference steps -0.5255 0.1844 0.4013 0.4285 0.4423 0.5135
Table 7: Performance (IR) comparison across different numbers of ensembled base models and varying inference steps.
Appendix H Comparison on Fewer Inference Steps
We compare our AFA with the baseline methods under fewer inference steps. As shown in Table 8, our AFA not only outperforms the baseline methods at higher inference steps but also demonstrates significantly better performance under fewer inference steps. This highlights AFA’s superior tolerance to the fewer inference steps compared to the baselines.
Group I Group II
inference steps 50 40 30 20 10 50 40 30 20 10
Wtd. Merging 0.3909 0.3104 0.2451 -0.0535 -0.7536 0.4387 0.4123 0.2942 0.1046 -0.5105
MBW 0.3922 0.3041 0.2345 0.0341 -0.7731 0.4396 0.4094 0.2893 -0.0031 -0.5524
autoMBW 0.3672 0.3158 0.2510 -0.4173 -0.8046 0.3513 0.3545 0.2502 -0.0841 -0.5841
MagicFusion 0.3317 0.3245 0.3104 0.2349 0.0175 0.4194 0.3995 0.3408 0.1951 0.0818
AFA (Ours) 0.4388 0.4389 0.4238 0.4191 0.3575 0.4892 0.4874 0.4846 0.4803 0.4013
Table 8: Quantitative comparison (IR) on fewer inference steps.
Appendix I Efficiency Comparison
We compare the efficiency of merging methods and ensembling methods. As shown in Table 9, when using the same 50 inference steps, ensembling methods (e.g., MagicFusion and AFA) do not have an efficiency advantage over a single base models or merging methods. The number of model parameters, TFLOPs, and inference time all increase linearly with the number of base models.
However, thanks to AFA’s high tolerance for fewer inference steps, it can achieve similar performance with reduced steps, as demonstrated in Table 7, while the baseline methods fail to maintain performance, as shown in Table 8. Therefore, under fewer inference steps, AFA achieves comparable efficiency (TFLOPs and inference time) to that of a single base models and merging methods.
2 base models 3 base models 6 base models
# params. TFLOPs Times (s) # params. TFLOPs Times (s) # params. TFLOPs Times (s)
Merging Method 859.52M 70.24 2.94 859.52M 70.24 2.94 859.52M 70.24 2.94
MagicFusion 1719.04M 138.66 6.01 2578.56M 206.88 9.13 5157.12M 411.53 22.41
AFA (50 inf. steps) 1761.47M 148.64 6.28 2621.03M 218.74 8.98 5210.37M 430.08 23.15
AFA (30 inf. steps) 1761.47M 94.24 3.21
AFA (20 inf. steps) 2621.03M 86.61 3.62
AFA (10 inf. steps) 5210.37M 103.68 4.58
Table 9: Efficiency Comparison between the merging methods and the ensembling methods when generating an image with a resolution of 512$\times$512. Note that only the parameters of denoisers are considered. Note that all the inference times are evaluated in our A6000 environment over 20 runs. Note that both the merging methods and MagicFusion are evaluated with 50 inference steps, because they cannot achieve comparable performance with fewer inference steps.
Appendix J Generality on Other Architectures
To evaluate the generality of our AFA on other diffusion architectures, we select three SDXL (Podell et al., 2023) models and two FLUX .1 [dev] [^16] https://blackforestlabs.ai/announcing-black-forest-labs/ models from CivitAI.
SDXL consists two U-Nets: the denoising U-Net and the refining U-Net. Both U-Nets shares a similar architecture of SDv1.5, but contain larger blocks. Consistent with the method used for ensembling SDv1.5 models, we employ the SABW module to aggregate features from each block.
FLUX .1 [dev] is a DiT-based model (Peebles & Xie, 2023). To ensemble models with this architecture, we also utilize the SABW module to aggregate features from each DiT Transformer block.
As shown in Table 10, when ensembling models with the SDXL or FLUX .1 [dev] architectures, the performance of the ensembled model surpasses that of the individual base models, demonstrating the generality of our AFA on different model architectures. However, the observed improvement is relatively small, likely because the base models already exhibit strong performance, and ensembling leads to only marginal gains. Additionally, due to the large size of the base models, ensembling multiple such models may be less practical.
It is worth noting that our AFA cannot be applied to ensembling models with different architectures, as the misalignment of block features makes block-wise aggregation challenging. We view ensembling models with diverse architectures as a promising direction for future research.
SDXL FLUX .1 [dev]
AES PS HPSv2 IR AES PS HPSv2 IR
Base Model A 6.1948 21.7854 28.0452 .6314 6.2341 21.8843 28.0958 .7341
Base Model B 6.1363 21.7783 28.0531 .6328 6.2346 21.8593 28.1003 .7324
Base Model C 6.1852 21.7752 27.9984 .6339
AFA (Ours) 6.2190 21.7959 28.0894 .6342 6.2490 21.8931 28.1194 .7370
Table 10: Quantitative comparison of AFA with different diffusion architectures.
Appendix K Distillation to Fewer Inference Steps
While our AFA demonstrates robustness to fewer inference steps, we aim to further distill the ensembled model into significantly fewer steps. Specifically, we explore using LCM (Luo et al., 2023) to achieve this. In our experiments, we kept the parameters of the denoisers frozen while training only the parameters of the SABW modules. However, as shown in Table 11, this approach is unsuccessful. A possible reason for this failure is that the frozen parameters may have impeded the distillation process.
     Original Ensembled Model      Distilled Ensembled Model
     Inference Steps      50      20      20      4      2      1
     IR      0.4892      0.4803      0.4793      -0.5846      -0.5594      -0.5952
Table 11: Experimental results (IR) for distilling the ensembled model with AFA into fewer inference steps using LCM.
Appendix L Evaluation on More Datasets
To validate the generality of our AFA, we evaluate it on additional datasets, which are DiffusionDB (Wang et al., 2022), JourneyDB (Pan et al., 2023), and LAION-COCO [^22] https://laion.ai/blog/laion-coco/, respectively. We randomly selected 50,000 samples from each dataset to compare the performance of AFA against the base models and the baseline methods [^23] Note that for JourneyDB, we ensured the selected samples are distinct from those in the training set..
The quantitative comparisons on the three additional datasets are presented in Table 12, Table 13, and Table 14, respectively. The results demonstrate that our AFA consistently outperforms both the base models and the baseline methods, highlighting the generality and effectiveness of our approach.
Group I Group II
FID $\downarrow$ IS CLIP-I CLIP-T AES PS HPSv2 IR FID $\downarrow$ IS CLIP-I CLIP-T AES PS HPSv2 IR
Base Model A 14.78 6.88 .6433 .2987 5.557 23.4113 29.9126 .4354 14.31 7.02 .6578 .3124 5.612 23.5527 30.1521 .4413
Base Model B 15.01 6.93 .6513 .3010 5.456 23.1123 29.8131 .4223 14.98 7.01 .6618 .3128 5.502 23.3128 30.0035 .4307
Base Model C 15.13 6.87 .6441 .2988 5.478 23.3423 29.9341 .4339 15.02 6.94 .6391 .2983 5.481 23.3420 29.9413 .4291
Wtd. Merging 17.41 6.37 .6214 .2776 5.213 23.2975 28.9938 .4007 16.98 6.81 .6231 .2984 5.501 23.2139 28.7310 .4147
MBW 16.67 6.81 .6378 .2843 5.367 23.3002 28.8393 .4115 15.20 6.99 .6392 .2931 5.493 23.4412 29.0012 .4216
autoMBW 16.66 6.76 .6342 .2663 5.377 23.3874 28.8432 .4293 15.17 7.00 .6381 .2891 5.551 23.3584 29.8413 .4298
MagicFusion 14.77 6.98 .6561 .3123 5.674 23.4132 29.9241 .4440 14.21 7.02 .6641 .3125 5.611 23.5049 29.9941 .4391
AFA (Ours) 12.32 7.39 .6765 .3691 5.773 23.5132 30.0946 .4763 13.98 7.11 .6712 .3211 5.698 23.6130 30.1931 .4712
Table 12: Quantitative comparison in Group I and II on DiffusionDB.
Group I Group II
FID $\downarrow$ IS CLIP-I CLIP-T AES PS HPSv2 IR FID $\downarrow$ IS CLIP-I CLIP-T AES PS HPSv2 IR
Base Model A 17.80 5.63 .5324 .2484 4.520 19.2066 24.5997 .3606 17.21 5.75 .5378 .2513 4.577 19.1089 24.5196 .3633
Base Model B 18.17 5.72 .5379 .2501 4.466 18.9965 24.5139 .3458 18.01 5.72 .5428 .2588 4.482 18.9069 24.3972 .3483
Base Model C 17.98 5.62 .5270 .2434 4.507 19.1876 24.5743 .3554 17.87 5.61 .5169 .2398 4.445 18.9523 24.2945 .3443
Wtd. Merging 20.98 5.26 .5170 .2314 4.288 19.1237 23.8558 .3327 20.55 5.50 .5108 .2452 4.455 18.8782 23.3782 .3399
MBW 19.74 5.58 .5218 .2365 4.399 19.1381 23.7068 .3371 18.35 5.74 .5221 .2378 4.484 19.0703 23.5806 .3465
autoMBW 20.07 5.54 .5204 .2227 4.429 19.1960 23.6664 .3508 18.04 5.70 .5182 .2325 4.501 18.9612 24.2498 .3510
MagicFusion 17.76 5.76 .5412 .2591 4.672 19.2147 24.5909 .3666 16.93 5.72 .5414 .2573 4.550 19.0955 24.3816 .3570
AFA (Ours) 14.86 6.04 .5587 .3063 4.763 19.3324 24.7316 .3849 16.72 5.78 .5482 .2646 4.651 19.1986 24.5480 .3990
Table 13: Quantitative comparison in Group I and II on JourneyDB.
Group I Group II
FID $\downarrow$ IS CLIP-I CLIP-T AES PS HPSv2 IR FID $\downarrow$ IS CLIP-I CLIP-T AES PS HPSv2 IR
Base Model A 13.82 7.82 .7328 .3432 6.260 26.5012 33.8765 .4962 13.26 8.06 .7549 .3559 6.402 26.8726 34.4362 .5080
Base Model B 14.12 7.89 .7397 .3467 6.164 26.1997 33.7818 .4754 13.87 8.05 .7583 .3617 6.286 26.5950 34.2635 .4910
Base Model C 14.08 7.73 .7280 .3371 6.197 26.4257 33.8805 .4886 13.84 7.90 .7284 .3379 6.254 26.6099 34.1549 .4879
Wtd. Merging 16.46 7.24 .7101 .3209 5.886 26.3613 32.8551 .4573 15.79 7.76 .7149 .3415 6.271 26.5130 32.8277 .4762
MBW 15.45 7.68 .7243 .3275 6.071 26.3787 32.6951 .4688 14.11 8.07 .7344 .3327 6.311 26.7870 33.1185 .4860
autoMBW 15.61 7.67 .7218 .3101 6.092 26.4808 32.6464 .4852 13.98 8.00 .7328 .3280 6.343 26.6503 34.0794 .4946
MagicFusion 13.80 7.92 .7477 .3594 6.425 26.5094 33.8799 .5048 13.13 8.05 .7632 .3632 6.421 26.8431 34.2797 .5024
AFA (Ours) 11.66 8.36 .7716 .4241 6.582 26.6673 34.1098 .5331 12.84 8.16 .7717 .3735 6.552 26.9677 34.4844 .5405
Table 14: Quantitative comparison in Group I and II on LAION-COCO.
Appendix M Ensembling Models with Highly Correlated Features
As the number of the base models increases, reliance on multiple base models may leads to overfitting to correlated features. To evaluate the robustness of our AFA against highly correlated features, we design an experiment. Specifically, we selected one high-quality base model and one low-quality base model. Using AFA, we ensemble the high-quality model with several low-quality models to evaluate whether the ensembled model will perform toward that of the low-quality model due to the dominance of the correlated features from the low-quality model.
As shown in Table 15, although the performance the ensembled model slightly declines when ensembling with large number of low-quality models, it still outperforms the high-quality model. This demonstrates that our AFA exhibits strong robustness to highly correlated features.
     AES      PS      HPSv2      IR
     High-Quality Base Model      5.5641      21.8013      28.0183      .4238
     Low-Quality Base Model      5.5013      21.4624      27.7246      .2835
     AFA (1 High-Quality Model + 1 Low-Quality Model)      5.6013      21.9341      28.3485      .4746
     AFA (1 High-Quality Model + 2 Low-Quality Models)      5.6000      21.9348      28.3398      .4739
     AFA (1 High-Quality Model + 3 Low-Quality Models)      5.5974      21.8248      28.3399      .4730
     AFA (1 High-Quality Model + 5 Low-Quality Models)      5.5831      21.8019      28.2035      .4593
Table 15: Quantitative comparison of ensembling models with high-correlated features.
Appendix N Training with Dataset of Lower Quality
In this paper, we train our AFA using a high-quality dataset, JourneyDB. This raises the concern of whether AFA’s high performance is primarily attributed to the quality of the training dataset. To address this, we also train AFA using LAION-COCO, a dataset of lower quality compared to JourneyDB.
As shown in Table 16, the model trained on LAION-COCO performs similarly to the one trained on JourneyDB, indicating that AFA is not sensitive to the quality of the training dataset. Its consistent performance across datasets of varying quality highlights its robustness and generalization ability. Regardless of whether the training dataset is high-quality or low-quality, AFA effectively leverages the features of the base models to deliver stable and reliable results.
Group I Group II
FID $\downarrow$ IS CLIP-I CLIP-T AES PS HPSv2 IR FID $\downarrow$ IS CLIP-I CLIP-T AES PS HPSv2 IR
JourneyDB 9.76 7.14 .6926 .2675 5.5201 21.8263 27.9734 .4388 10.27 7.42 .6855 .2717 5.5798 21.8059 28.0371 .4892
LAION-COCO 9.74 7.16 .6930 .2664 5.5093 21.8257 27.9781 .4390 10.29 7.39 .6861 .2711 5.5789 21.8077 28.0363 .4889
Table 16: Quantitative comparison in Group I and II trained on JourneyDB and LAION-COCO.
Appendix O More Experimental Details
We train our AFA in an environment equipped with 8 NVIDIA V100 GPUs, each with 32GB of memory. When ensembling two base models, the GPU memory consumption is about 11GB under FP16 precision and a batch size of 1. When ensembling three base models, the memory consumption increases to 14GB under the same settings. For six base models, the memory consumption reaches about 30GB. Due to the frozen parameters of the base models, training our AFA demands relatively modest GPU memory, even when ensembling a larger number of base models.